OpenStudy (anonymous):

Find the exact value of sin(cos^-1(3/7)) is the answer 2sqrt10/7 or 7/2sqrt10

OpenStudy (anonymous):

A

OpenStudy (anonymous):

2sqrt(10)/7

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ x^2+3^2=7^2\implies x=\sqrt{49-9}=\sqrt{40} = 2\sqrt{10} \implies x/7 = 2\sqrt{10}/7 \]

OpenStudy (anonymous):

thanks to both!I get it now